11. The expected average rate of return for a proposed investment of $800,000 in a fixed asset with a useful life of 4 years, straight-line depreciation, no residual value, and an expected total net income of $360,000 for the 4 years, is a. 45% b. 22.5% c. 11.3% d. 5.5% the answer is b but why out
Answer #1

Average investment = 800000 / 2 = 400000

Average income per year = 360000/4=90000

ROI = 90000 / 400000 = .225 OR 22.5%

Correct option is "B"